OpenStudy (anonymous):

2x-z=10 y+2z=-2 3x-2y=8 please help

OpenStudy (anonymous):

2x-z=10 -> Z = ....?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2x-10?

OpenStudy (anonymous):

Correct, now replace 2x-10 into y+2z=-2

OpenStudy (anonymous):

y+4x-20

OpenStudy (anonymous):

nm

OpenStudy (anonymous):

y + 2z = -2 y + 4x -20 = -2 Can you simplify it?

OpenStudy (anonymous):

y+4x=18

OpenStudy (anonymous):

3x-2y = 8 4x + y = 18 Can you solve it?

OpenStudy (anonymous):

7x-y=26?

OpenStudy (anonymous):

No, you need to multiply 2 ( 4x + y = 18 )

OpenStudy (across):

\[\begin{bmatrix} 2 & 0 & -1\\ 0 & 1 & 2\\ 3 & -2 & 0 \end{bmatrix}\begin{bmatrix} x\\ y\\ z \end{bmatrix}=\begin{bmatrix} 10\\ -2\\ 8 \end{bmatrix}\Longrightarrow\begin{bmatrix} x\\ y\\ z \end{bmatrix}=\begin{bmatrix} 4\\ 2\\ -2 \end{bmatrix}.\]Use that to check your solution.